OpenStudy (anonymous):

Jace must read a chapter of his book for a homework assignment. The chapter has thirty-two full pages and Jace has already read seventeen point five . Which equation will tell us the number of pages Jace still has to read

OpenStudy (anonymous):

Do you have answer choices? owo

OpenStudy (anonymous):

yes a) p+32=17.5 b) 17.5+32=p c) 32+p=17.5 d) p+17.5=32

OpenStudy (anonymous):

If you were to express that paragraph with a wordy equation, you'd get something like Pages Left = Total Pages - Pages Read So just substitute the numbers in and you'll get P = 32 - 17.5 That can then be rearranged to reach an answer that matches your given choices

OpenStudy (anonymous):

It'd be D because p+17.5=32 Subtract 17.5 on both sides! So...: 32-17.5=p And 32-17.5 is 14.5 So p=14.5

OpenStudy (anonymous):

thankyou so much!